### v2

@@ Abstract
This work is motivated by (i) the recently proposed relationship between gravitational tension (through the Kretschmann scalar) and energy density Estrada and Aros (2025) to construct regular black hole (RBH) solutions in Pure Lovelock (PL) gravity, and (ii) the fact that, for PL, the inclusion of a negative cosmological constant $\Lambda<0$ leads to the appearance of a potential curvature singularity Cai and Ohta (2006). Thus, by using a particular choice of coupling constants such that the resulting equations of motion for Lovelock gravity have an $n$-fold degenerate ground state (LnFDGS) AdS, we provide a methodology to construct RBH solutions with $\Lambda<0$ using an energy density model analogous to (but different in the definition of gravitational tension from) the one mentioned earlier. Moreover, because relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution is complicated, we define an alternative version of the Kretschmann scalar suitable for the LnFDGS AdS theory and, consequently, a redefinition of the gravitational tension. Remarkably, we obtain a model where there exists a value $r_{*}$ slightly greater than the extremal radius, $r_{*}>r_{ext}$, which could be on the order of the Planck length, such that the solutions of the vacuum AdS black hole and our AdS RBH become indistinguishable. However, at short length scales such that $r<r_{*}$, quantum effects would arise, causing both cases to differ in their geometry (suppressing the central singularity) and their thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior.
@@ I Introduction
Regarding the aforementioned points, in this work, we will provide a methodology to study the analogy between gravitational tension and energy density in order to construct regular black hole solutions in LnFDGS gravity under the presence of a negative cosmological constant. It is worth mentioning that directly relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution in the same way as for PL theory is complicated, because several additional terms would appear in the gravitational tension. Due to this, unlike Estrada and Aros (2025), we will define an alternative version of the Kretschmann scalar suitable for the LnFDGS AdS theory, and consequently, a redefinition of the gravitational tension. Additionally, we will speculatively discuss the connection between our model and the quantum Schwinger effect for this theory. We will test how the presence of the negative cosmological constant influences the obtained RBH solution, both in terms of its geometric and thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we will propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior.
@@ VI.2.1 On the radial evolution
On the other hand, related to the analysis of phase transitions, this can be easily unfolded by using the fact that $C=\left(\frac{dT}{dM}\right)^{-1}\sim\left(\frac{dT}{dr_{+}}\right)^{-1}$ in graph 11. Moving from right to left, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature having to rise to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. As mentioned earlier, this could occur at scales close to the Planck scale. That is, the nature of our matter sources, along with avoiding the formation of a singularity, would lead to the formation of a remnant. Thus, in our solution, there would be two phase transitions, from right to left: i) $C>0$ where $dT/dM>0$, ii) $C<0$ where $dT/dM<0$ (these two coincide with the vacuum LnFDGS AdS case, where both solutions are indistinguishable), iii) $C>0$ where $dT/dM>0$.
@@ VII DISCUSSION AND SUMMARIZE
Regarding the analysis of heat capacity and radial evolution, moving from right to left in graph 11, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature rising to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. We could assume that at Planck scales, our matter sources induce a phase transition such that the black hole becomes stable at these scales, leading to a black hole remnant rather than a divergent temperature.

### v3

@@ Abstract
In Estrada and Aros (2025), a relationship between gravitational tension (GT) and energy density, via the Kretschmann scalar (KS), was recently proposed to construct regular black hole (RBH) solutions in Pure Lovelock (PL) gravity. However, in PL gravity it is not viable to include a negative cosmological constant, as this leads to the appearance of a potential curvature singularity Cai and Ohta (2006). In this work, by choosing a particular set of coupling constants such that the resulting equations of motion for Lovelock gravity admit an $n$-fold degenerate ground state (LnFDGS) AdS solution, we construct an RBH solution with $\Lambda<0$, providing an energy density model analogous to (but distinct in its definition of GT) the one previously mentioned. Moreover, since relating the gravitational tension to the KS of the vacuum LnFDGS solution is nontrivial, we provide an alternative definition of both the KS and the GT. Remarkably, we obtain a model where there exists a value $r_{*}$ slightly greater than the extremal radius, $r_{*}>r_{ext}$, which could be on the order of the Planck length, such that the solutions of the vacuum AdS black hole and our AdS RBH become indistinguishable. However, at short length scales such that $r<r_{*}$, quantum effects would arise, causing both cases to differ in their geometry (suppressing the central singularity) and their thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior. We also provide a speculative methodology to compare theoretical results from Lovelock gravity with experimental results from the EHT for M87. Graphically, we have found that, depending on the chosen parameters, the theoretical results obtained for Lovelock gravity can match the experimental data.
@@ I Introduction
Regarding the aforementioned points, in this work, we will provide a methodology to study the analogy between gravitational tension and energy density in order to construct regular black hole solutions in LnFDGS gravity under the presence of a negative cosmological constant. It is worth mentioning that directly relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution in the same way as for PL theory is complicated, because several additional terms would appear in the gravitational tension. Due to this, unlike Estrada and Aros (2025), we will define an alternative version of the Kretschmann scalar suitable for the LnFDGS AdS theory, and consequently, a redefinition of the gravitational tension. Additionally, we will speculatively discuss the connection between our model and the quantum Schwinger effect for this theory. We will test how the presence of the negative cosmological constant influences the obtained RBH solution, both in terms of its geometric and thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we will propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior. We also provide a speculative way to compare theoretical results obtained for Lovelock gravity with the experimental results from the EHT for M87.
@@ V Shadows
V.2 A brief and speculative glimpse into the M87-EHT constraints
Here, we aim to provide a speculative glimpse into the behavior of the shadow sizes obtained in the previous subsection, in relation to the constraints captured by the EHT, concerning the shadow size of the four-dimensional supermassive black hole M87. Within the $1\sigma$ (68%) confidence level, it can be found that the shadow size of the aforementioned black hole lies within the interval Nozari et al. (2024); Cao et al. (2023); Kocherlakota et al. (2021):
$4.31\leq\frac{2}{V_{\text{vac}}(r=r_{\text{sh}})}=\frac{r_{\text{sh}}}{M}\leq 6.08$ (50)
where $V_{\text{vac}}(r)$ corresponds to the absolute value of the gravitational potential of the vacuum case:
$V_{\text{vac}}(r)=\frac{2M}{r}$ (51)
[…]
$V^{(LL)}_{\text{vac}}(r)=\left(\frac{2M}{r^{d-2n-1}}\right)^{d-2n-1}$ (52)
In this way, and in a speculative manner, we write the ratio of constraint (50) as follows:
$4.31\leq\frac{2^{d-2n-1}}{V^{(LL)}_{\text{vac}}(r=r_{\text{sh}})}=\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}\leq 6.08$ (53)
so, it is direct to check that the ratio $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ is dimensionless and for $n=1,d=4$ the constraint (53) reduces to eq (50).
In Figures 11 and 12, we display the behavior of $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ versus the horizontal axis, which corresponds to a fraction of $M$. In Figure 11, we consider the case $n=2$ with $d=6,7,8$, and in Figure 12, the case $n=3$ with $d=8,9,10$. We observe that, depending on the chosen parameters, the behavior of the vacuum gravitational potential $\frac{2^{d-2n-1}}{V^{(LL)}{\text{vac}}(r=r{\text{sh}})}$ is such that the obtained results can be adjusted to those reported by the EHT for M87 (within the context of four dimensions and General Relativity). In this way, we provide a speculative methodology to compare theoretical results from Lovelock gravity with experimental results from the EHT for M87. Particularly, we find that our results match the experimental results from the EHT for M87 in our plots corresponding to $n=2,\ d=7,8$ and $n=3,\ d=10,11$. It appears that, for a fixed value of $n$, the parameters tend to better fit the experimental results as the number of dimensions increases.
Figure 11: In vertical axis $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ for $n=2,a=4,l=100$. First panel: On the horizontal axis $M/10$ for $d=6$. Second panel: On the horizontal axis $M/100$ for $d=7$. Third panel: On the horizontal axis $M/1000$ for $d=8$
Figure 12: In vertical axis $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ for $n=3,a=4,l=100$. First panel: On the horizontal axis $M/100$ for $d=9$. Second panel: On the horizontal axis $M/1000$ for $d=10$. Third panel: On the horizontal axis $M/10000$ for $d=11$
@@ VI.2.1 On the radial evolution
On the other hand, related to the analysis of phase transitions, this can be easily unfolded by using the fact that $C=\left(\frac{dT}{dM}\right)^{-1}\sim\left(\frac{dT}{dr_{+}}\right)^{-1}$ in graph 13. Moving from right to left, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature having to rise to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. As mentioned earlier, this could occur at scales close to the Planck scale. That is, the nature of our matter sources, along with avoiding the formation of a singularity, would lead to the formation of a remnant. Thus, in our solution, there would be two phase transitions, from right to left: i) $C>0$ where $dT/dM>0$, ii) $C<0$ where $dT/dM<0$ (these two coincide with the vacuum LnFDGS AdS case, where both solutions are indistinguishable), iii) $C>0$ where $dT/dM>0$.
@@ VII DISCUSSION AND SUMMARIZE
We have also provided a speculative way to compare theoretical results obtained for Lovelock gravity with the experimental results from the EHT for M87. To this end, we have used the vacuum gravitational potential of the Lovelock solution as a comparison parameter, such that our Equation (53) reduces to Constraint (50) for $n=1,\ d=4$, which corresponds to the data provided by the EHT for M87. We tested whether the vacuum gravitational potential yields values consistent with the constraints given by Equation (53). Graphically, we have found that, depending on the chosen parameters, the theoretical results obtained for Lovelock gravity can match the experimental data. In particular, we find that our results agree with the experimental results from the EHT for M87 in our plots corresponding to $n=2,\ d=7,8$ and $n=3,\ d=10,11$. It seems that, for a fixed value of $n$, the parameters tend to better match the experimental results as the number of dimensions increases.
@@
Regarding the analysis of heat capacity and radial evolution, moving from right to left in graph 13, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature rising to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. We could assume that at Planck scales, our matter sources induce a phase transition such that the black hole becomes stable at these scales, leading to a black hole remnant rather than a divergent temperature.